using this information. The marginal private benefits for a good is represented by MPB = 80 - Q while the marginal social benefits for the good is MSB = 89-Q. Additionally the marginal cost of consuming good A is represented by the equation MC = 20 + 2Q.

A. What is the level of consumption that will occur if only the private benefits are considered?

B. What is the efficient amount?
C. Is it more or less than the amount found if only private benefits are considered.
D. This is an example of a positive or negative externality
Answer #1

A. If only the private benefits are considered, the equilibrium condition is:

MPB = MC

80 - Q = 20 + 2Q

3Q = 60

Q = 60 / 3 = 20

Thus, the level of consumption is 20 units.

B. The efficient amount is at the point where,

MSB = MC

89 - Q = 20 + 2Q

3Q = 69

Q = 69 / 3 = 23

Thus, the efficient amount is 23 units.

C. The efficient amount is more than the amount found if only private benefits are considered.

D. This is an example of a positive externality, because the efficient amount is more than the amount found if only private benefits are considered.